OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). sin θ cos 3θ + cos θ sin 3θ = 0

OpenStudy (anonymous):

sin u cos v +cos u sin v = sin (u+v)

OpenStudy (anonymous):

θ=(nπ)/4

OpenStudy (anonymous):

\[\sin 4\theta= 0 /theta= pi k/4

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

is that all the solutions?

OpenStudy (anonymous):

no,0,(π)/4,(π)/2,(3π)/4,π,(5π)/4,(3π)/2,(7π)/4

OpenStudy (anonymous):

I forgot 0